# On the convergence of noisy Bayesian Optimization with Expected Improvement

### v1

@@ Abstract
Expected improvement (EI) is one of the most widely-used acquisition functions in Bayesian optimization (BO). Despite its proven success in applications for decades, important open questions remain on the theoretical convergence behaviors and rates for EI. In this paper, we contribute to the convergence theories of EI in three novel and critical area. First, we consider objective functions that are under the Gaussian process (GP) prior assumption, whereas existing works mostly focus on functions in the reproducing kernel Hilbert space (RKHS). Second, we establish the first asymptotic error bound and its corresponding rate for GP-EI with noisy observations under the GP prior assumption. Third, by investigating the exploration and exploitation of the non-convex EI function, we prove improved error bounds for both the noise-free and noisy cases. The improved noiseless bound is extended to the RKHS assumption as well.
@@
Bayesian optimization, expected improvement, convergence rate, Gaussian process prior, noisy observations, non-convex, exploration, exploitation
@@ 1 Introduction
Bayesian optimization (BO) is a derivative-free optimization method for black-box functions [1, 2, 3]. BO has seen enormous success in many applications such as machine learning [4], structural design [5], robotics [6], additive manufacturing [7], inertial confinement fusion design [8], etc. In the classic form, BO aims to solve the optimization problem
@@
A surrogate model, often by default a Gaussian process (GP), is used to approximate the black-box objective function (1)  [9]. An acquisition function guides the selection of sequential points for observations. One of the most successful and widely used acquisition functions is the expected improvement (EI) [10]. EI calculates the conditional expectation of an improvement function so that both the mean value and variance of the GP are used in search of the next sample. With a closed form, it is simple to implement in that only the cumulative distribution function (CDF) and probability density function (PDF) of the standard normal distribution are required. Thanks to its effectiveness and efficiency, EI has witnessed rich extensions in literature, e.g., scalable methods [11], constrained EI [12], etc. The classic noise-free BO algorithm using EI is also referred to as the efficient global optimization (EGO) algorithm [13].
@@
Despite its wide adoption and success, existing works on the theoretical convergence of Bayesian optimization with expected improvement, hereby referred to as GP-EI throughout the paper, have clear gaps. The two main streams of research study the asymptotic convergence and the cumulative regret bound of BO algorithms. The asymptotic convergence analysis of GP-EI is limited by the assumptions on the objective and whether observation noise is allowed. Under the assumptions that $f$ is in the reproducing kernel Hilbert space (RKHS) of the GP kernel, also called the frequentist setting, and no observation noise, the asymptotic convergence rates of GP-EI is established in [14], where the convergence rates differ based on the kernels used in the GP. Specifically, define the error $r_{t}^{0}:=f({\bm{x}}_{t}^{+})-f^{*}$, where ${\bm{x}}_{t}^{+}$ is the sample that produces the best current observation among $t$ samples and $f^{*}=\underset{\begin{subarray}{c}{\bm{x}}\in C\end{subarray}}{\text{argmin}}f({\bm{x}})$. The error bound is of $\mathcal{O}(t^{-\frac{1}{d}})$ for squared exponential (SE) kernels and $\mathcal{O}(t^{-\frac{\min\{\nu,1\}}{d}}\log^{\alpha}(t))$ for Matérn kernels, where $\nu$ is the parameter of the kernel and $\alpha=\frac{1}{2}$ for $\nu\in\mathbb{N}$ in Matérn kernels and $\alpha=0$ otherwise. However, to the best of our knowledge, the extensions of [14] to noisy GP-EI or when $f$ follows different assumptions, e.g., $f$ is sampled from a GP, are yet to be accomplished. Other works related to asymptotic convergence of GP-EI include [15], where the authors used no-empty-ball properties of kernels to show that the sequence of iterates generated by GP-EI are dense. In [16], the author showed that for the ranking and selection problems, a modified EI is essentially identical to optimal computing budget allocation (OCBA) algorithm, via the asymptotic sampling ratios. In [17], the author proved consistency of EI for noiseless GP with continuous sample paths.
@@
The other stream of research has focused on the cumulative regret bound, based on the seminal work of [18]. In [19], the authors added additional steps and auxiliary hyper-parameters to modify EI. Further, they assumed bounded conditions in RKHS on the hyper-parameters to derive a cumulative regret bound. In [20], the authors used a stopping criterion to help bound the prediction error. However, the analysis relies on a critical yet not established result that the standard deviation at the optimal point decreases at the same rate as sample points. In [21], the authors attempted to bound the standard deviation at the optimal point via structured initial sampling and achieved an improvement on the regret bound for SE kernels. While effective, their method requires an infinite number of initial samples asymptotically. In [22], the authors proved no-regret for their modified EI, which includes additional control parameters for EI that could become unbounded as $t$ increases.
@@
In this paper, we address gaps in the asymptotic convergence analysis of GP-EI mentioned above. Our contributions can be summarized as follows. First, we prove the first asymptotic convergence rate for noisy GP-EI under the GP prior assumption for $f$ (see Assumption 2.4). Given $\delta\in(0,1)$, we prove the upper bound for the noisy error measure $r_{t}:=y^{+}_{t}-f^{*}$, where $y^{+}_{t}$ is the best current observation among $t$ samples, is $\mathcal{O}\left(t^{-\frac{1}{2}}\log(t)^{\frac{d+1}{2}}\right)$ for SE kernels and $\mathcal{O}(t^{\frac{-\nu}{2\nu+d}}\log^{\frac{\nu}{2\nu+d}}(t))$ for Matérn kernels, with probability greater than $1-\delta$. In the noisy GP-EI case, the convergence rate refers to the rate of decrease of the upper bound of $r_{t}$, or equivalently, the rate at which $f({\bm{x}}_{t}^{+})$ converges to the vicinity of the $f^{*}$. Second, we present the asymptotic convergence rates without noise under the GP prior assumption, extending the results in [14] which are obtained under the RKHS assumption. Third, we prove improved error bounds for GP-EI with or without noise under the GP prior assumption. The improved error bounds stem from our novel quantification of the trade-off between exploitation and exploration of EI, the latter being reflected in the posterior standard deviation, at non-sample points. The emerging analysis technique is general and can be used to analyze other convergence measure, e.g., the regret of GP-EI. As an example, we prove a better error bound compared to [14], under the RKHS assumption and noiseless case.
This paper is organized as follows. In Section 2, we describe the basis of the GP-EI algorithm and other background information. In Section 3, we prove multiple properties of EI and the preliminary theoretical results under GP prior assumptions, which are the foundation of the analysis in Section 4. We present our asymptotic convergence analysis on noisy GP-EI under GP prior assumptions in Section 4. In Section 4.4, we improve asymptotic convergence results to noiseless GP-EI under RKHS assumptions. Numerical illustrations of the theories are presented in Section 5, while conclusions are offered in Section 6.
@@ 2.1 Gaussian process
where ${\bm{k}}_{t}({\bm{x}})=[k({\bm{x}}_{1},{\bm{x}}),\dots,k({\bm{x}}_{t},{\bm{x}})]^{T}$. When there is no observation noise, (2) can still be used to compute $\mu_{t}$ and $\sigma_{t}$ with a small $\sigma$, e.g., $\sigma=10^{-8}$, for stability of matrix inversion [27, 28].
@@ 2.2 Expected improvement
where
(5) $z_{t}({\bm{x}})=\frac{y^{+}_{t}-\mu_{t}({\bm{x}})}{\sigma_{t}({\bm{x}})}.$
The functions $\phi$ and $\Phi$ are the PDF and CDF of the standard normal distribution, respectively.
@@
From the assumptions on kernels in Section 2.1, the exploration $b$ is constrained by $b\in(0,1]$. We view $a$ and $b$ as two independent variables. For a given ${\bm{x}}$, if $a_{t}=y^{+}_{t}-\mu_{t}({\bm{x}})$ and $b_{t}=\sigma_{t}({\bm{x}})\in[0,1]$, we have $EI(a_{t},b_{t})=EI_{t}({\bm{x}})$.
Another commonly used function [21] in the analysis of EI is $\tau:\mathbb{R}\to\mathbb{R}$ defined as
@@
In order to solve (8), optimization algorithms including L-BFGS or random search can be used. The GP-EI algorithm is given in Algorithm 1, where a stopping criterion can be added.
@@ 2.3 Additional background
We present the remaining necessary background information and our assumptions in this section. In order to find the rate of decrease of the error bound in the noisy case, we use the following well-established lemma for GP based on information theory [18].
@@
The sum of GP posterior variance $\sigma_{t}({\bm{x}}_{t+1})$ at next sample ${\bm{x}}_{t+1}$ satisfies
@@
For Gaussian process with $t$ samples, the SE kernel has
$\gamma_{t}=\mathcal{O}(\log^{d+1}(t))$, and the Matérn kernel with smoothness parameter $\nu>0$ has $\gamma_{t}=\mathcal{O}(t^{\frac{d}{2\nu+d}}(\log^{\frac{2\nu}{2\nu+d}}(t)))$.
@@
Consider a positive definite kernel $k:C\times C\to\mathbb{R}$ with respect to a finite Borel measure supported on $C$. A Hilbert space $H_{k}$ of functions on $C$ with an inner product $\langle\cdot,\cdot\rangle_{H_{k}}$ is called a RKHS with kernel $k$ if for $\forall{\bm{x}}\in C$, $k(\cdot,{\bm{x}})\in H_{k}$, and $\forall{\bm{x}}\in C,f\in H_{k}$, $\langle f,k(\cdot,{\bm{x}})\rangle_{H_{k}}=f({\bm{x}})$. The induced RKHS norm $\left\lVert{f}\right\rVert_{H_{k}}=\sqrt{\langle f,f\rangle_{H_{k}}}$ measures the smoothness of $f$ with respect to $k$.
For ease of reference, we list the assumptions used in our analysis below, all of which are common in BO literature. First, we have the GP prior assumption for $f$, also called the Bayesian setting in literature.
@@
The objective function $f$ is in the RKHS $\mathcal{H}_{k}$ associated with the kernel $k({\bm{x}},{\bm{x}}^{\prime})\leq 1$ and $k({\bm{x}},{\bm{x}})=1$. Further, the RKHS norm of $f$ is bounded $\left\lVert{f}\right\rVert_{\mathcal{H}_{k}}\leq B$ for some $B\geq 1$.
@@ 3.1 EI properties
The last statement in Lemma 3.1 is a well-known result (e.g., see proof of Lemma 5.1 in [18]), as Figure 1 (left) demonstrates. The property of $\tau(\cdot)$ in (7) is given below.
@@
From the definition of $\tau(z)$, we can write
@@
Thus, the derivative of $\tau$ is
(12) $\frac{d\tau(z)}{dz}=\Phi(z)+z\phi(z)-\phi(z)z=\Phi(z)>0,$
which completes the proof.
@@
Thus, if $z>\frac{1+\sqrt{5}}{2}$, then the right-hand-side of (15) is greater than $\phi(-z)$ and
(16) $q(z):=\Phi(-z)-\tau(-z)=(1+z)\Phi(-z)-\phi(-z)>0.$
@@
For $z>1$, $\frac{d^{2}q(z)}{dz^{2}}>0$. For $z\in(0,1)$ $\frac{d^{2}q(z)}{dz^{2}}<0$. Thus, $q^{\prime}(z)$ is monotonically increasing for $z>1$ and decreasing for $z\in(0,1)$. For $z>1$, from Lemma 3.2,
@@
Further, $q^{\prime}(0)=\Phi(0)-\phi(0)>0$ and $q^{\prime}(1)=\Phi(-1)-\phi(-1)<0$. Thus, there exists unique $\bar{z}\in(0,1)$ so that $q^{\prime}(\bar{z})=0$. For $z\in(0,\bar{z})$, $q(z)$ is monotonically increasing. For $z\in[\bar{z},\frac{1+\sqrt{5}}{2}]$, $q(z)$ is monotonically decreasing. Therefore, $q(z)>\min\{q(0),q(\frac{1+\sqrt{5}}{2})\}$ for $z\in(0,\frac{1+\sqrt{5}}{2})$. Since $q(0)>0$ and $q(\frac{1+\sqrt{5}}{2})>0$, $q(z)>0$ for $z\in(0,\frac{1+\sqrt{5}}{2})$. Combined with (16), the proof is complete.
@@
If $z_{t}({\bm{x}})<0$, or equivalently $y^{+}_{t}-\mu_{t}({\bm{x}})<0$, (21) leads to $\frac{EI_{t}({\bm{x}})}{\sigma_{t}({\bm{x}})}<\phi(z_{t}({\bm{x}}))$. If $z_{t}({\bm{x}})\geq 0$, $\Phi(\cdot)<1$ gives us $\frac{EI_{t}({\bm{x}})}{\sigma_{t}({\bm{x}})}<z_{t}({\bm{x}})+\phi(z_{t}({\bm{x}}))$. The left inequality in (20) is a direct result of $EI_{t}({\bm{x}})\geq y^{+}_{t}-\mu_{t}({\bm{x}})$.
The monotonicity of the exploration and exploitation of (6) is given next.

### v2

# On the convergence rate of noisy Bayesian Optimization with Expected Improvement

@@ Abstract
Expected improvement (EI) is one of the most widely used acquisition functions in Bayesian optimization (BO). Despite its proven success in applications for decades, important open questions remain on the theoretical convergence behaviors and rates for EI. In this paper, we contribute to the convergence theory of EI in three novel and critical areas. First, we consider objective functions that fit under the Gaussian process (GP) prior assumption, whereas existing works mostly focus on functions in the reproducing kernel Hilbert space (RKHS). Second, we establish for the first time the asymptotic error bound and its corresponding rate for GP-EI with noisy observations under the GP prior assumption. Third, by investigating the exploration and exploitation properties of the non-convex EI function, we establish improved error bounds of GP-EI for both the noise-free and noisy cases.
@@
Bayesian optimization, global optimization, expected improvement, convergence rate, Gaussian process, noisy observations, non-convex, exploration, exploitation
@@ 1 Introduction
Bayesian optimization (BO) is a derivative-free optimization method for black-box functions [1, 2, 3] that enjoys enormous success in many applications such as hyperparameter optimization in machine learning [4], structural design [5], robotics [6], additive manufacturing [7], inertial confinement fusion design [8], etc. In recent years, BO has been increasingly applied to novel area such as expensive integral objectives [9], distributionally robust optimization [10], risk-averse optimization [11], etc. In the classic form, BO aims to solve the optimization problem
@@
A surrogate model, often a Gaussian process (GP), is used to approximate the black-box objective function (1)  [12]. An acquisition function guides the selection of sequential points for observations. One of the most successful and widely used acquisition functions is the expected improvement (EI) [13]. EI calculates the conditional expectation of an improvement function such that both the mean value and variance of the GP are used in search of the next sample. With a closed form, it is simple to implement in that only the cumulative distribution function (CDF) and probability density function (PDF) of the standard normal distribution are required. Thanks to its effectiveness and efficiency, EI has witnessed rich extensions in literature, e.g., scalable methods [14], constrained EI [15], etc. The classic noise-free BO algorithm using EI is also referred to as the efficient global optimization (EGO) algorithm [16].
@@
Despite its wide adoption and success, existing works on the theoretical convergence of Bayesian optimization with expected improvement, hereby referred to as GP-EI throughout the paper, have clear gaps. The two main streams of research study the asymptotic convergence and the cumulative regret bound of BO algorithms, the latter focusing on the performance over the entire optimization path. The asymptotic convergence analysis of GP-EI is limited by the assumptions on the objective and whether observation noise is allowed. Under the assumptions that $f$ is in the RKHS of the GP kernel, also called the frequentist setting, with no observation noise, the asymptotic convergence rates of GP-EI is established in [17], where the convergence rates differ based on the kernels used in the GP. Specifically, define the error $r_{t}^{0}:=f({\bm{x}}_{t}^{+})-f^{*}$, where ${\bm{x}}_{t}^{+}$ is the sample that produces the best current observation among $t$ samples and $f^{*}=\underset{\begin{subarray}{c}{\bm{x}}\in C\end{subarray}}{\text{argmin}}f({\bm{x}})$. The error bound is $\mathcal{O}(t^{-\frac{1}{d}})$ for squared exponential (SE) kernels and $\mathcal{O}(t^{-\frac{\min\{\nu,1\}}{d}}\log^{\alpha}(t))$ for Matérn kernels, where $\nu$ is the parameter of the Matérn kernel and $\alpha=\frac{1}{2}$ for $\nu\in\mathbb{N}$ and $\alpha=0$ otherwise. However, to the best of our knowledge, the extensions of [17] to noisy GP-EI or when $f$ follows different assumptions, e.g., $f$ is sampled from a GP, are yet to be established.
@@
The other stream of research has focused on the cumulative regret bound, based on the seminal work of [18]. In [19], the authors added additional steps and auxiliary hyper-parameters to modify EI. Further, they assumed bounded conditions in RKHS on the hyper-parameters to derive a cumulative regret bound. In [20], the authors used a stopping criterion to help bound the prediction error. However, the analysis relies on a critical yet not established result that the standard deviation at the optimal point decreases at the same rate as the sampled points. In [21], the authors attempted to bound the standard deviation at the optimal point via structured initial sampling and achieved an improvement on the regret bound for SE kernels. Their method requires an infinite number of initial samples as the total number of samples increases. In [22], the authors proved no-regret for their modified EI, which includes additional control parameters for EI that could become unbounded as $t$ increases.
@@
Other works related to asymptotic convergence of GP-EI include [28], where the authors showed that the sequence of iterates generated by GP-EI are dense In [29], the author showed that for the ranking and selection problems, a modified EI is identical to optimal computing budget allocation (OCBA) algorithm, via the asymptotic sampling ratios. In [27], the author proved consistency of EI for noiseless GP with continuous sample paths.
In this paper, we address gaps in the asymptotic convergence analysis of GP-EI mentioned above. Our contributions can be summarized as follows. First, we prove for the first time the asymptotic convergence rate for noisy GP-EI under the GP prior assumption for $f$ (see Assumption 2.5). Given $\delta\in(0,1)$, we prove that the upper bound for the noisy error measure $r_{t}:=y^{+}_{t}-f^{*}$, where $y^{+}_{t}$ is the best current observation among $t$ samples, is $\mathcal{O}\left(t^{-\frac{1}{2}}\log(t)^{\frac{d+1}{2}}\right)$ for SE kernels and $\mathcal{O}(t^{\frac{-\nu}{2\nu+d}}\log^{\frac{\nu}{2\nu+d}}(t))$ for Matérn kernels, with probability greater than $1-\delta$. In the noisy GP-EI case, the convergence rate refers to the rate of decrease for $r_{t}$, or equivalently, the rate at which $y_{t}^{+}$ reduces to values no larger than $f^{*}$. Second, we present the asymptotic convergence rates without noise under the GP prior assumption, thereby extending the results in [17], which are obtained under the RKHS assumption. Notably, we achieve the same convergence rates as [17] when using the same kernels. Third, we prove improved error bounds for GP-EI with or without noise under the GP prior assumption. The improved error bounds stem from our novel quantification of the trade-off between exploitation and exploration of EI, the latter being reflected in the posterior standard deviation, at non-sample points. The emerging analysis technique is general and can be used to analyze other convergence measure, e.g., the cumulative regret of GP-EI. As an example, we prove a better error bound compared to [17], under the RKHS assumption and noiseless case.
This paper is organized as follows. In Section 2, we describe the basis of the GP-EI algorithm and other background information. In Section 3, we prove multiple properties of EI and the preliminary theoretical results under GP prior assumptions, which are the foundation of the analysis in Section 4. We present our asymptotic convergence analysis on noisy GP-EI under GP prior assumptions in Section 4. In Section 4.3, we improve asymptotic convergence results to noiseless GP-EI under RKHS assumptions. Numerical illustrations of our theoretical developments are presented in Section 5, while conclusions are offered in Section 6.
@@ 2.1 Gaussian process
where ${\bm{k}}_{t}({\bm{x}})=[k({\bm{x}}_{1},{\bm{x}}),\dots,k({\bm{x}}_{t},{\bm{x}})]^{T}$.
@@ 2.2 Expected improvement
where $z_{t}({\bm{x}})=\frac{y^{+}_{t}-\mu_{t}({\bm{x}})}{\sigma_{t}({\bm{x}})}$. The functions $\phi$ and $\Phi$ are the PDF and CDF of the standard normal distribution, respectively.
@@
From the assumptions on kernels in Section 2.1, the exploration parameter $b$ is constrained by $b\in(0,1]$. We view $a$ and $b$ as two independent variables. For a given ${\bm{x}}$, if $a_{t}=y^{+}_{t}-\mu_{t}({\bm{x}})$ and $b_{t}=\sigma_{t}({\bm{x}})\in[0,1]$, $EI(a_{t},b_{t})=EI_{t}({\bm{x}})$.
Another commonly used function in the analysis of EI is $\tau:\mathbb{R}\to\mathbb{R}$ defined as
@@
In order to solve (7), optimization algorithms such as L-BFGS or random search can be used. The GP-EI algorithm is given in Algorithm 1, where a stopping criterion can be used, e.g., a prescribed computational budget.
@@ 2.3 Additional background
We present the remaining necessary background information and our assumptions in this section. Our analysis uses repeatedly union bound, or Boole’s inequality, which we provide in the following lemma for completeness.
@@
For a countable set of events $A_{1},A_{2},\dots$, we have
$\mathbb{P}(\bigcup_{i=1}^{\infty}A_{i})\leq\sum_{i=1}^{\infty}\mathbb{P}(A_{i}).$
In order to find the rate of decrease of the error bound in the noisy case, we use the following well-established lemma for GP based on information theory [18].
Lemma 2.2.
The sum of GP posterior variance $\sigma_{t}$ at next sample ${\bm{x}}_{t+1}$ satisfies
@@
For a GP with $t$ samples, the SE kernel has $\gamma_{t}=\mathcal{O}(\log^{d+1}(t))$, and the Matérn kernel with smoothness parameter $\nu>0$ has $\gamma_{t}=\mathcal{O}(t^{\frac{d}{2\nu+d}}(\log^{\frac{2\nu}{2\nu+d}}(t)))$.
@@
Consider a positive definite kernel $k:C\times C\to\mathbb{R}$ with respect to a finite Borel measure supported on $C$. A Hilbert space $H_{k}$ of functions on $C$ with an inner product $\langle\cdot,\cdot\rangle_{H_{k}}$ is called a RKHS with kernel $k$ if $k(\cdot,{\bm{x}})\in H_{k}$ for all ${\bm{x}}\in C$, and $\langle f,k(\cdot,{\bm{x}})\rangle_{H_{k}}=f({\bm{x}})$ for all ${\bm{x}}\in C,f\in H_{k}$. The induced RKHS norm $\left\lVert{f}\right\rVert_{H_{k}}=\sqrt{\langle f,f\rangle_{H_{k}}}$ measures the smoothness of $f$ with respect to $k$.
For ease of reference, we list the assumptions used in our analysis below, all of which are common in literature. First, we have the GP prior assumption for $f$, also called the Bayesian setting. We further assume that $f$ is Lipschitz continuous, as in [18].
@@
The objective function $f$ is in the RKHS $\mathcal{H}_{k}$ associated with the kernel $k({\bm{x}},{\bm{x}}^{\prime})\leq 1$ and $k({\bm{x}},{\bm{x}})=1$. Further, the RKHS norm of $f$ is bounded $\left\lVert{f}\right\rVert_{\mathcal{H}_{k}}\leq B$ for some $B\geq 1$. Moreover, $f$ is Lipschitz continuous with constant $L$.
@@ 3.1 EI properties
The last statement in Lemma 3.1 is a well-known result (e.g., see proof of Lemma 5.1 in [18]). Figure 1 (left) also illustrates this point. The property of $\tau(\cdot)$ in (6) is given below.
@@
From the definition (6) of $\tau(z)$, we can write
@@
Thus, direct differentiation of $\tau(z)$ gives
$\frac{d\tau(z)}{dz}=\Phi(z)+z\phi(z)-\phi(z)z=\Phi(z)>0.$
@@
Thus, if $z>\frac{1+\sqrt{5}}{2}$, then the right-hand side of (12) is greater than $\phi(-z)$ and
(13) $q(z)=\Phi(-z)-\tau(-z)=(1+z)\Phi(-z)-\phi(-z)>0.$
@@
For $z>1$, $\frac{d^{2}q(z)}{dz^{2}}>0$. For $z\in(0,1)$ $\frac{d^{2}q(z)}{dz^{2}}<0$. Thus, $q^{\prime}(z)$ is monotonically increasing for $z>1$ and decreasing for $z\in(0,1)$. For $z>1$, from Lemma 3.2 and definition of $\tau$,
@@
Further, $q^{\prime}(0)=\Phi(0)-\phi(0)>0$ and $q^{\prime}(1)=\Phi(-1)-\phi(-1)<0$. Thus, there exists unique $\bar{z}\in(0,1)$ so that $q^{\prime}(\bar{z})=0$. For $z\in(0,\bar{z})$, $q(z)$ is monotonically increasing. For $z\in[\bar{z},\frac{1+\sqrt{5}}{2}]$, $q(z)$ is monotonically decreasing. Therefore, $q(z)>\min\{q(0),q(\frac{1+\sqrt{5}}{2})\}$ for $z\in(0,\frac{1+\sqrt{5}}{2})$. Since $q(0)>0$ and $q(\frac{1+\sqrt{5}}{2})>0$, we have $q(z)>0$ for $z\in(0,\frac{1+\sqrt{5}}{2})$. Combined with (13), the proof is complete.
@@
If $z_{t}({\bm{x}})<0$, or equivalently $y^{+}_{t}-\mu_{t}({\bm{x}})<0$, (18) leads to $\frac{EI_{t}({\bm{x}})}{\sigma_{t}({\bm{x}})}<\phi(z_{t}({\bm{x}}))$. If $z_{t}({\bm{x}})\geq 0$, $\Phi(\cdot)<1$ gives us $\frac{EI_{t}({\bm{x}})}{\sigma_{t}({\bm{x}})}<z_{t}({\bm{x}})+\phi(z_{t}({\bm{x}}))$. The left inequality in (17) is a direct consequence of $EI_{t}({\bm{x}})\geq y^{+}_{t}-\mu_{t}({\bm{x}})$.
The monotonicity of the exploration and exploitation of form (5) is given next.